# Extracting the Speed of Sound in Heavy-Ion Collisions: A Study of Quantum Fluctuations and Thermalization

### v2

@@ Abstract
Probe of thermalization.– While the exact value of the speed of sound is of great significance to QCD dynamics, we emphasize that the process of extracting the speed of sound also serves as a direct probe of the thermalization of the quark-gluon plasma (QGP), which is also important. The response between mean transverse momentum and particle yield fundamentally differs in a locally thermalized system compared to a non-thermal system. These differences can be effectively distinguished by analyzing the quantum fluctuations of $\delta$ on an event-by-event basis.
@@
Based on our findings, we propose a combined measurement of the speed of sound and the Gaussianity of $\delta$ to characterize local thermalization. In a thermalized system, one expects the extracted speed of sound to have a physical value that falls within a reasonable range Cherman et al. (2009); Hippert et al. (2025), despite uncertainties such as those arising from the determination of the effective temperature, $T_{\rm eff}$, which is partially model-dependent. Simultaneously, the probability distribution $\rho(\delta|c_{s}^{2})$ should exhibit Gaussian behavior, up to corrections of order $1/N_{0}^{a}$. In contrast, if thermalization is not achieved, neither a physical value of the speed of sound nor Gaussianity in $\delta$ should be observed.
@@
Moreover, deviations from thermalization can be quantified using higher-order cumulants of the $\rho(\delta|c_{s}^{2})$ distribution. For example, the standardized kurtosis, defined as $\kappa_{4}=\{\delta^{4}\}/\{\delta^{2}\}^{2}-3$, should vanish in a system with complete local thermalization but remain finite in the presence of non-thermal contributions. From our numerical simulations, we find $\kappa_{4}\gtrsim 2.0$ in both HIJING and PYTHIA models. In contrast, hybrid hydrodynamic modeling yields $\kappa_{4}\approx 0.16$ for p-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV and $8.16$ TeV, and $\kappa_{4}\approx-0.06$ for Pb-Pb collisions, indicating small but finite non-thermal contributions.
@@
Summary and discussion. – For a QGP system created in ultracentral nucleus-nucleus collisions, if it is locally thermalized, the response relation between $\Delta_{p}$ and $\Delta_{N}$ is thermodynamic. Accordingly, event-by-event fluctuations, which are primarily quantum in origin, should be independent of the thermodynamic response and follow a Gaussian distribution. Although these fluctuations are correlated with $\Delta_{N}$ and induce corrections to averages, the Gaussianity condition of $\delta$ allows for the statistical extraction of the physical value of the speed of sound, for example, by solving the equation derived from the condition $\{\delta^{3}\}_{c}=0$. We expect the coefficients in this equation to be experimentally measurable, as they correspond to the standardized skewness or mixed skewness of the joint probability distribution $\mathcal{P}(\Delta_{N},\Delta_{p})$. Therefore, the physical value of the speed of sound can be precisely determined in experiments, even in systems where fluctuations are substantial, such as in small colliding systems like p-Pb.
@@
The non-Gaussianity of $\delta$ can serve as a robust diagnostic tool for thermalization in QGP systems. In realistic heavy-ion collisions, especially in smaller systems, local thermalization is expected to be partial, with non-thermal contributions becoming more significant in non-central collisions. As a result, we predict a monotonic increase in $\kappa_{4}$ of $\delta$ as a function of centrality. When $\kappa_{4}$ of $\delta$ exceeds unity, the dominant contribution to the response between $\Delta_{p}$ and $\Delta_{N}$ is no longer thermodynamic, indicating that the system should not be considered thermalized. The same argument extends to systems with finite baryon density, where the characterization of thermalization becomes even more crucial for the search for the QCD critical point.
@@
Acknowledgements. – We are grateful to Jean-Yves Ollitrault, Wei Li, and Jiangyong Jia for very helpful discussions. We also thank Xin-Nian Wang for suggesting the HIJING model and Xiangyu Wu for the HIJING code support. Y.-S. M. would like to thank Chen Zhong and Wen-Hao Zhou for their assistance with the computation server. This work is supported in part by the National Natural Science Foundation of China through Grants No. 12375133 and No. 12147101 (L.Y.), No. 12147101, No. 12225502, and No. 12075061 (X.-G. H.), by the Natural Science Foundation of Shanghai through Grant No. 23JC1400200, and by the National Key Research and Development Program of China through Grant No. 2022YFA1604900. The computations are performed at the CFFF platform of Fudan University.
Supplemental Material
The fluid-like fireball created in heavy-ion collisions converts to particles on a hyper-surface $\Sigma$, subject to the freeze-out condition. Through this conversion, the energy and entropy of the expanding fireball are expressed as
[…]
where $\sigma^{\mu}$ denotes the normal vector to the freeze-out hyper-surface, and $T^{\mu\nu}$ and $s^{\mu}$ are the relativistic energy-momentum tensor and entropy flow vector of the fluid, respectively. With respect to the locally equilibrated equation of state (LEOS) that relates energy or entropy density to temperature, the fireball can be effectively reduced to a uniform system characterized by the effective temperature $T_{\rm eff}$ and effective volume $V_{\rm eff}$, as described by the second equation in Eq. (6).
Particles generated from the freeze-out hyper-surface evolve further according to the UrQMD model. The resulting mean transverse momentum $\langle p_{T}\rangle$ of these final particles is expected to exhibit a linear correlation with the effective temperature $T_{\rm eff}$ on an event-by-event basis. For Pb-Pb collisions, which are approximately 2+1 dimensional due to longitudinal boost invariance, we find that $\langle p_{T}\rangle^{(i)}\approx 3T_{\rm eff}^{(i)}$, where $i$ labels the event.
In the p-Pb system, however, longitudinal boost invariance no longer holds, introducing ambiguity in the total energy determined by Eq. (6). The effective temperatures obtained using the $(t,x,y,z)$ coordinate system and the $(\tau,x,y,\eta_{s})$ Milne coordinate system—where $\tau=\sqrt{t^{2}-z^{2}}$ and $\eta_{s}=\tanh^{-1}(z/t)$—are found to differ. Despite this discrepancy, solutions from both coordinate systems exhibit a strong linear correlation with $p_{T}^{(i)}$ on an event-by-event basis, although the correlation coefficients differ. This difference implies that, in a fireball without boost invariance, the mean transverse momentum of the final-state particles may receive contributions from the kinematic energy associated with longitudinal expansion. As there are two sets of linear relations to the effective temperatures, we consider the realistic energy scale of the fireball that is characterized by the mean transverse momentum in the extraction of the speed of sound lies in between, which gives rise to a range in the linear relation between $\langle p_{T}\rangle^{(i)}$ and $T_{\rm eff}^{(i)}$.

### v3

@@ Abstract
Probe of thermalization.– While the exact value of the speed of sound is of great significance to QCD dynamics, we emphasize that the process of extracting the speed of sound also serves as a direct probe of the thermalization of the QGP, which is also important. The response between mean transverse momentum and particle yield fundamentally differs in a locally thermalized system compared to a non-thermal system. These differences can be effectively distinguished by analyzing the statistical properties of $\delta$ on an event-by-event basis.
@@
Based on our findings, we propose a combined measurement of the speed of sound and the Gaussianity of $\delta$ to characterize local thermalization. In a thermalized system, one expects the extracted speed of sound to have a physical value that falls within a reasonable range Cherman et al. (2009); Hippert et al. (2025), despite uncertainties such as those arising from the determination of the effective temperature, $T_{\rm eff}$, which is partially model-dependent. Simultaneously, the probability distribution $\rho(\delta|c_{s}^{2})$ should exhibit Gaussianity, up to corrections of order $1/N_{0}^{a}$. In contrast, if thermalization is not achieved, neither a physical value of the speed of sound nor Gaussianity in $\delta$ should be observed.
@@
Moreover, deviations from thermalization can be quantified using higher-order cumulants of the $\rho(\delta|c_{s}^{2})$ distribution. For example, the standardized kurtosis, defined as $\kappa_{4}=\{\delta^{4}\}/\{\delta^{2}\}^{2}-3$, should vanish in a system with complete local thermalization but remain finite in the presence of non-thermal contributions. From our numerical simulations, we find $\kappa_{4}=2.15\pm 0.05$ and $1.91\pm 0.11$ in HIJING and PYTHIA models, respectively. In contrast, hybrid hydrodynamic modeling yields $\kappa_{4}=0.16\pm 0.05$ and $0.18\pm 0.07$ for p-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV and $8.16$ TeV, and $\kappa_{4}=-0.064\pm 0.096$ for Pb-Pb collisions, signaling small but finite non-thermal contributions.
@@
Summary and discussion. – For a QGP system created in ultracentral nucleus-nucleus collisions, if it is locally thermalized, the response relation between $\Delta_{p}$ and $\Delta_{N}$ is thermodynamic. Accordingly, event-by-event fluctuations, which are primarily quantum in origin, should be independent of the thermodynamic response and follow a Gaussian distribution. Although these fluctuations are correlated with $\Delta_{N}$ and induce corrections to averages, the Gaussianity condition of $\delta$ allows for the statistical extraction of the physical value of the speed of sound, for example, by solving the equation derived from the condition $\{\delta^{3}\}_{c}=0$. We expect the coefficients in this equation to be experimentally measurable, as they correspond to the standardized skewness or mixed skewness of the joint probability distribution $\mathcal{P}(\Delta_{N},\Delta_{p})$. Therefore, the physical value of the speed of sound can be determined in experiments, even in systems where fluctuations are substantial, such as in small colliding systems like p-Pb. Note however, the inferred speed of sound is subject to uncertainties arising from initial-state modeling and viscous corrections hyd , which warrant further systematic investigation through dedicated theoretical studies.
@@
The non-Gaussianity of $\delta$ can serve as a robust diagnostic tool for thermalization in QGP systems. When $\kappa_{4}$ of $\delta$ exceeds unity, the dominant contribution to the response between $\Delta_{p}$ and $\Delta_{N}$ is no longer thermodynamic, indicating that the system should not be considered thermalized. Noting that a deterministic thermodynamic response involving $c_{s}^{2}$ remains valid in the presence of finite baryon density, the current formulation is expected applicable at lower collision energies nex , where the characterization of thermalization becomes even more crucial for the search of the QCD critical point.
@@
Acknowledgements. – We are grateful to Jean-Yves Ollitrault, Wei Li, and Jiangyong Jia for very helpful discussions. We also thank Xin-Nian Wang for suggesting the HIJING model and Xiangyu Wu for the HIJING code support. Y.-S. M. would like to thank Chen Zhong, Lu-Meng Liu and Wen-Hao Zhou for their assistance with the computation server. This work is supported in part by the National Natural Science Foundation of China through Grants No. 12375133 and No. 12147101 (L.Y.), No. 12147101, No. 12225502, and No. 12075061 (X.-G. H.), by the Natural Science Foundation of Shanghai through Grant No. 23JC1400200, and by the National Key Research and Development Program of China through Grant No. 2022YFA1604900. The computations are performed at the CFFF platform of Fudan University.